OpenStudy (anonymous):

Which set of polar coordinates describes the same location as the rectangular coordinates (4,-4)? (Points : 2)

OpenStudy (anonymous):

theta = arctan(-4/4) = 7*pi/4 r= sqrt(4^2 +(-4)^2)